OpenStudy (anonymous):

WILL MEDAL AND FAN!!!!!!! Complete the equation for the linear function whose graph contains the point (7, 8) and has a slope of −1.2.

OpenStudy (anonymous):

@Walkerman15

Directrix (directrix):

y - 8 = -1.2* ( x - 7) y - 8 = -1.2x + 8.4 y = -1.2x + 16.4 ---------> check this @OGangsta

OpenStudy (anonymous):

thank you